# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ 2.2 Data Annotation
To validate our models and enable further qualitative analysis, we build on the human-annotated dataset introduced in kostikova-fine (57), which comprises 2,864 instances (1,437 for migrants, 1,427 for women) from the DeuParl corpus. [^1] Annotation was carried out by five trained student assistants through an iterative process, achieving Cohen’s $\kappa$ of 0.62 (high level) and 0.42 (fine-grained). Final labels in the dataset are assigned by majority vote. Each instance is labeled according to a fine-grained solidarity framing scheme that distinguishes four high-level stance categories – solidarity (willingness to share resources, following Lahusen.2018 (60) and TwitterDataset (50)), anti-solidarity (restriction or exclusion of groups), mixed (both supportive and opposing elements), and none (neutral or unrelated content). Instances labeled as solidarity or anti-solidarity are further assigned one of four frame types adapted from thijssen2022s (98): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance in contributions), and empathic (respect for or rejection of social diversity).
We extend the evaluation to a newly annotated set of 756 instances for Migrant with more historical coverage (before 1950). For this set, annotations were conducted by 4 annotators following the same procedure as in kostikova-fine (57) Statistics on the human-annotated dataset are presented in Fig. 3. Overall, the label distribution differs notably by target group: woman-related instances are predominantly labeled as solidarity (74.1%), while migrant-related instances show a more diverse pattern, with lower solidarity (51.6%) and a higher share of anti-solidarity (21.3%).
Figure 3: Distribution of instances in the human annotated dataset across time and target groups. Total indicates number of all annotated samples for the given target group, including the None label. See Table 2 for the actual numbers of instances in the human annotated dataset.
@@ 2.3 Models and Experiments
We evaluate a diverse suite of models, ranging from smaller supervised models (BERT, fine-tuned SBERT) to proprietary LLMs (GPT-4) and recent open-weight instruction-tuned models. The final list includes:
@@
    smaller pre-trained LMs like BERT and SBERT;
@@
    closed-source LLMs like gpt-4-1106-preview;
@@
    open-weight LLMs such as Llama-3-8B-It, Llama-3.3-70B-It (and fine-tuned Llama-3-70B-It), Gemma-2-9b-It, Gemma-2-27b-It, and Mistral-2-Large.
@@
We adopt the 70/15/15 train/dev/test split strategy as in kostikova-fine (57), generating three random splits of approximately 2,000 training, 400 development, and 430 test instances for each target group. These are used to train and evaluate supervised models. For inference-only experiments, we use only the test sets: Test 1 (original) and Test 2 (newly added, 756 instances), the latter offering broader historical coverage (1860s–1950s). We report Macro F1 to account for class imbalance, along with per-class F1 scores for both high-level and fine-grained classification tasks. [^2] High-level classification refers to the four stance categories: solidarity, anti-solidarity, mixed, and none. Fine-grained classification includes ten labels: the four subtypes of solidarity, the four subtypes of anti-solidarity, plus mixed and none.
Models
@@
We use manually designed prompts for each target group. Each prompt consists of a classification instruction concatenated with the parliamentary instance, i.e., the keyword sentence plus its three-sentence left/right context as defined in 2.1. Note that, although speaker and party metadata are available in the corpus, this information is removed from the model input. The classification instruction includes (i) chain-of-thought reasoning in zero- and few-shot settings (via think step by step); (ii) label definitions; and (iii) a two-step classification format: high-level first, followed by subtype for solidarity or anti-solidarity prediction. [^3] For GPT-4, both steps are combined in a single prompt (see Figure 21 for Migrant and Figure 22 in the Appendix for Woman). For open-weight models, we instead use two prompts: one for high-level classification and, conditional on that output, a second for subtype selection (see Figure 24 for Migrant and Figure 27 in the Appendix for Woman). In few-shot mode, prompts include one representative example per label.
We apply these prompts to closed-source models like gpt-4-1106-preview, and to open-weight LLMs including Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it (full precision) in inference-only setting. Larger models (Llama-3.3-70B-Instruct and Mistral-Large-Instruct-2407) are run with $Q4_{0}$ quantization. Additionally, we fine-tune Llama-3-8B-It and Llama-3-70B-It using a three-model pipeline (high-level and two subtype classifiers), applying instruction-style fine-tuning with true labels paired with GPT-4-generated explanations as training targets.
We use BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details for all models used in our experiments are provided in C.
@@ 3 Results
In this section, we evaluate models against a human-annotated dataset to determine reliability for social scientific analysis of solidarity frames in parliamentary texts. Our aim is twofold: (i) to identify a model whose predictions approximate human-level annotation quality across both high-level and fine-grained stance categories, and (ii) to ensure the selected model is practically feasible to deploy at scale. Specifically, we ask whether recent open-weight models can match the performance of GPT-4, which was previously found to be highly reliable (though costly) kostikova-fine (57), through techniques such as few-shot prompting or lightweight fine-tuning.
3.1 Model Performance Overview
Table 1 reports macro F1 scores for each model across both inference-based and fine-tuned approaches. We benchmark these results against a human upper bound, estimated using a leave-one-out (LOO) procedure, [^4] For each annotator, we exclude their labels, compute a consensus label from the remaining annotators, and calculate the macro F1 score between the excluded annotator and the consensus. We then average these scores across all annotators. and report performance for both high-level and fine-grained classification tasks, with the latter shown in parentheses. Our observations are structured around two dimensions: (i) overall model performance, (ii) task and label-related challenges.
@@ 3.1 Model Performance Overview
Few-shot 0.37 (0.54) 0.43 (0.63) 0.29 (0.51)
@@
Few-shot 0.30 (0.50) 0.37 (0.65) 0.28 (0.53)
Llama-3-70B Fine-tuned 0.26 (0.49) 0.27 (0.65) 0.21 (0.47)
Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.17 (0.40)
Few-shot 0.14 (0.35) 0.19 (0.34) 0.15 (0.40)
Fine-tuned 0.16 (0.38) 0.20 (0.40) 0.13 (0.27)
Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.25 (0.45)
Few-shot 0.25 (0.55) 0.33 (0.48) 0.26 (0.46)
Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.22 (0.46)
Few-shot 0.23 (0.44) 0.31 (0.54) 0.24 (0.46)
BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.23 (0.45)
SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.18 (0.36)
Human upper bound 0.34 (0.62) 0.48 (0.75) 0.35 (0.57)
@@
Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks. ?Migrant (T1)? uses the original benchmark; ?Migrant (T2)? is newly introduced in this paper. Green cells indicate the best-performing score for each target group, based on a model’s best setting (e.g., zero-shot or few-shot); yellow cells indicate the best-performing score from a different model (i.e., the second-best model overall).
  1. (i)
    Overall model performance.
    • GPT-4 achieves the highest scores, but Llama-3.3-70B is a strong alternative. Across all target groups (Woman and both Migrant test sets), GPT-4 (zero-shot) consistently shows the top performance, coming closest to the human upper bound in both high-level and fine-grained classification. Llama-3.3-70B is the best-performing open model, showing meaningful gains from few-shot prompting (in particular in the high-level classification).
@@
    • Model size correlates with performance, but not perfectly. Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure 4 for the relationship between model size and performance). Mid-sized models follow, and smaller models perform worst. However, this relationship is not strictly linear: for example, Llama3-8B-it underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B-it.
@@
      Figure 4: Model size vs. performance on high-level and fine-grained classification tasks in Migrant (T1). Each model is plotted with its best-performing configuration (zero- or few-shot). The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
    • Examples help mid-sized models more than small or very large ones. Few-shot prompting improves performance primarily for mid-sized models such as Gemma-9B-it, Gemma-27B-it, Mistral-2-Large, and Llama-3.3-70B-it. In contrast, smaller models like Llama-3-8B-it show no benefit, and GPT-4 sees minimal gains—likely because its zero-shot performance is already near-optimal. Fine-tuned variants (e.g., Llama-3.3-70B-
      it
      , Llama-3-8B-it) often show no improvement over inference-only models.
  2. (ii)
    Task and label-related challenges.
    • All models perform worse on the fine-grained task, underscoring its complexity: unlike the more intuitive high-level categories (e.g., solidarity, anti-solidarity), it requires alignment with annotation guidelines grounded in social theory, where even human annotators show limited agreement (reaching macro F1 score 0.34 for Woman and 0.48 for Migrant).
    • At the high level, solidarity is the easiest to detect (see Figure 5; full scores in Table 3 in the Appendix), especially in the Woman category, where even smaller models like BERT (0.77) and Gemma-9B (0.78) perform strongly. Anti-solidarity and mixed stance are the most challenging categories, with GPT-4 achieving the highest high-level anti-solidarity scores (0.64 for Migrant, 0.65 for Woman), but mid-size models (e.g., Mistral, Llama-3.3-70B, Gemma-27B, Gemma-9B) show moderate performance (with scores ranging from 0.52 to 0.54). Only GPT-4 and Llama-3.3-70B reliably handle mixed stances (e.g., 0.38–0.43 for Migrant, 0.30–0.37 for Woman).
    • On the fine-grained level, GPT-4, Llama-3.3-70B, and Gemma models offer the most balanced coverage across subtypes. GPT-4 stands out on exchange-based solidarity (0.56), group- and exchange-based anti-solidarity for Migrant (0.47 and 0.46), and on empathic and compassionate anti-solidarity for Woman (0.67 and 0.50). Among open models, Llama-3.3-70B performs particularly well on compassionate solidarity (0.51), and shows strong results on group- and exchange-based anti-solidarity (0.42 and 0.35). Interestingly, the Gemma models are able to distinguish the rare compassionate anti-solidarity subtype for Migrant with some success. Most other models struggle with fine-grained anti-solidarity, with scores close to zero.

### v2

@@ 2.2 Data Annotation
To validate our models and enable qualitative analysis, we use a human-annotated dataset created in two batches. The first, introduced in Kostikova et al. (2024), comprises 2,864 instances (1,428 migrant-related, 1,437 woman-related) from DeuParl, annotated by five trained student assistants. The annotation guidelines were refined over multiple rounds of discussion among the annotators and authors; details of the annotation process and the final guidelines are provided in Appendix˜F and Appendix˜G in the Appendix. The second adds 756 migrant-related instances with broader historical coverage (pre-1950), annotated by four of the same assistants following the same procedure.
Each instance is labeled according to a fine-grained solidarity framing scheme distinguishing four high-level stance categories – solidarity, defined as willingness to share resources Lahusen and Grasso (2018); Ils et al. (2021), anti-solidarity (restriction or exclusion), mixed (both supportive and opposing elements), and none (neutral or unrelated content). Solidarity and anti-solidarity are further assigned one of four subtypes adapted from Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance), and empathic (respect for or rejection of diversity). We refer to the subtypes as fine-grained annotation.
The final gold-standard label for each instance is determined by majority vote.For the first batch, inter-annotator agreement reached a Cohen’s $\kappa$ of 0.62 for the high-level labels and 0.42 for the fine-grained labels. For the second batch, agreement reached 0.47 for the high-level labels and 0.37 for the fine-grained labels. We attribute the lower agreement in the second batch to the inclusion of older historical data, which we discuss further in Section˜3.1. Statistics on the human annotated dataset are shown in Figure˜3. Label distributions differ by target group: woman-related instances are predominantly solidarity (74.1%), while migrant-related instances are more diverse, with less solidarity (51.6%) and more anti-solidarity (21.3%).
Figure 3: Distribution of human annotated instances by time and target group. Total indicates the number of instances per group (including None); see Table˜6 in the Appendix for exact counts.
@@ 2.3 Models and Experiments
We evaluate a diverse set of models to cover variation in model size, provider, and access setting (open vs. closed), while also including supervised classifier baselines. Our selection comprises:
@@
    smaller pre-trained LMs such as BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details are provided in Appendix˜C in the Appendix;
@@
    closed-source LLMs such as gpt-4-1106-preview and gpt-5-chat;
@@
    open-weight LLMs, including:
    • gpt-oss-20B and gpt-oss-120B;
    • Mistral-2-Large (123b parameters);
    • Llama-4-Scout-17B-16E-Instruct;
    • Llama-3.3-70B-It, Llama-3-8B-It;
    • Qwen2.5-72B-Instruct;
    • Gemma-2-9B-It and Gemma-2-27B-It.
@@
We adopt a 70/15/15 train/dev/test split of the first annotated batch (Kostikova et al., 2024), generating three random splits of about 2,000 training, 400 development, and 430 test (Test 1) instances in total, covering both target groups. Supervised models are trained and evaluated on these splits, while inference-only experiments use two test sets: Test 1 (original) and Test 2 (756 newly added migrant-related instances, 1860s–1950s). We report Macro F1 to account for class imbalance, along with per-class F1 for high-level (four stance categories) and fine-grained (ten labels) classification. [^2] High-level: solidarity, anti-solidarity, mixed, none. Fine-grained: four solidarity subtypes, four anti-solidarity subtypes, plus mixed and none.
Prompting Setup
@@
For the inference-only experiments with LLMs, We use manually designed prompts for each target group. Each prompt consists of a classification instruction concatenated with the parliamentary instance (the keyword sentence plus three sentences of left/right context as defined in Section˜2.1). Speaker and party metadata are available in the corpus but removed from the model input. Instructions include (i) chain-of-thought cues in zero-/few-shot settings (think step by step); (ii) label definitions; and (iii) a two-step format: high-level classification followed by subtype prediction. [^3] For GPT-4, both steps are combined in a single prompt (Figure 25 and Figure 26 in the Appendix); for open-weight models, we use two prompts: one for high-level classification and, conditional on that, a second for subtype selection (Figure 27 and Figure 30 in the Appendix). In few-shot mode, one representative example per label is included.
Model Ensemble
We additionally evaluate a model ensemble to test whether combining strong open-weight LLMs improves performance relative to individual models. The ensemble setup also resembles the human annotation procedure in that final labels are derived from multiple judgments rather than a single annotator. The ensemble combines the best-performing open models, each in its strongest configuration (Table˜1), using majority vote. The label with the most votes is selected as the final prediction; in tie cases, votes are weighted by each model’s macro-F1 score as measured on the validation set. We report ensemble results in Section˜3.1.3.
Design-based Supervised Learning (DSL)
When employing LLMs as annotators, it may be that single labels are systematically preferred compared to human annotators, yielding biased estimates and incorrect conclusions. We use Design-based Supervised Learning (DSL; Egami et al. 2023) to correct such systematic biases. DSL combines large-scale surrogate outputs with a subsample of human-coded data to construct bias-corrected pseudo-outcomes for inference. In our application, the surrogate outputs are soft label distributions derived from a committee of LLMs.
For each parliamentary speech $i$, we observe surrogate labels $Q_{i}$ and covariates $X_{i}$. For a (small) subset, we also have human annotations $Y_{i}$. The surrogate labels $Q_{i}$ are soft-label distributions over annotation categories, computed as the proportion of committee models assigning each label, while the human labels $Y_{i}$ are distributions over categories constructed from multiple human annotators. Human annotations are observed only for a subset of instances under a decade-stratified annotation design. We assign each instance a decade-specific inclusion probability $\pi_{i}$, computed as the proportion of instances in that decade that received human annotation, and treat these as fixed design probabilities in the DSL correction.
DSL corrects downstream estimands by constructing a bias-adjusted pseudo-outcome
$\tilde{Y}_{i}=\hat{g}(Q_{i},X_{i})+\frac{R_{i}}{\pi_{i}}\big(Y_{i}-\hat{g}(Q_{i},X_{i})\big),$ (1)
where $R_{i}$ indicates whether instance $i$ is human-annotated, and $\hat{g}(Q_{i},X_{i})$ estimates the conditional expectation of human labels given surrogate outputs and covariates. We model $\hat{g}$ using a multinomial regression trained on soft targets and implement $K$-fold cross-fitting.
Because outcomes are multi-class, DSL is applied in a vector-valued form, giving a vector of DSL-adjusted class scores $\tilde{Y}_{i}$ for each instance. We use them directly in downstream analyses.
We estimate separate DSL models for the high-level and fine-grained tasks to avoid carrying subtype-level error from fine-grained predictions into aggregated high-level estimates.
@@ 3 Results
In this section, we evaluate the models in two steps. We first assess instance-level annotation performance against human labels, including overall performance, task-specific challenges, and systematic error patterns in Section˜3.1. We then examine whether these model outputs support reliable downstream inference by comparing hard-label, soft-label, and DSL-corrected trend estimates against human-derived reference trends in Section˜3.2.
3.1 Annotation Performance at the Instance Level
Table˜1 reports macro F1 scores for each model under inference and fine-tuned settings. We benchmark against the leave-one-out F1 score of human annotators, [^4] For each annotator, we exclude their labels, compute a consensus from the others, and calculate the macro F1 score between the excluded annotator and the consensus; scores are then averaged across annotators. and show results for both high-level and fine-grained tasks (the latter in parentheses). We evaluate instance-level annotation performance along four dimensions: (i) overall performance of individual models; (ii) task- and label-related challenges; (iii) model ensemble performance; and (iv) error analysis.
@@ 3.1 Annotation Performance at the Instance Level
Few-shot 0.37 (0.54) 0.43 (0.63) 0.26 (0.55)
GPT-5 Zero-shot 0.36 (0.53) 0.42 (0.67) 0.30 (0.60)
[…]
Large open-weight models
gpt-oss-20B Zero-shot 0.25 (0.43) 0.27 (0.55) 0.15 (0.49)
Few-shot 0.28 (0.47) 0.40 (0.56) 0.22 (0.49)
gpt-oss-120B Zero-shot 0.28 (0.55) 0.38 (0.66) 0.25 (0.55)
[…]
@@
Llama-4-Scout Zero-shot 0.25 (0.46) 0.28 (0.52) 0.20 (0.50)
Few-shot 0.25 (0.46) 0.41 (0.63) 0.25 (0.50)
Qwen-2.5 Zero-shot 0.21 (0.40) 0.37 (0.62) 0.23 (0.52)
Few-shot 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
@@
[…]
Mid- and small-scale open-weight models
Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.15 (0.40)
Few-shot 0.14 (0.35) 0.19 (0.34) 0.12 (0.42)
Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.21 (0.46)
Few-shot 0.25 (0.55) 0.33 (0.48) 0.21 (0.47)
Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.21 (0.47)
Few-shot 0.23 (0.44) 0.31 (0.54) 0.23 (0.47)
Classical baselines
BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.22 (0.48)
SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.17 (0.37)
Human reference 0.32 (0.67) 0.48 (0.78) 0.30 (0.58)
@@
Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks. ?Migrant (T1)? is the original benchmark; ?Migrant (T2)? is newly introduced in this paper. Green cells mark the best score per target group (model’s best setting); yellow cells mark the best score among open models (second best overall). Gray text indicates suboptimal settings for a given model.
3.1.1 Overall model performance
  • GPT-5 leads, but some open models offer a strong alternative. Across Woman and both Migrant test sets, GPT-5 (few-shot) matches human annotators on the high-level task and surpasses them on the fine-grained task. Among open models, gpt-oss-120B matches the human reference value in fine-grained performance and comes close to GPT-4 and GPT-5 at the high level.
    Among the remaining open-weight models, Llama-4-Scout, Qwen-2.5, and Llama-3.3-70B show the strongest performance. These models achieve comparable results across targets and evaluation levels (e.g., on Migrant T1, fine-grained F1 scores range from 0.38 to 0.43 and high-level scores from 0.59 to 0.65). While their relative strengths vary by setting, where Qwen-2.5 and Llama-4-Scout achieve the strongest fine-grained scores, and Llama-3.3-70B the highest high-level scores, their overall performance remains closely clustered.
@@ 3.1.1 Overall model performance
  • Model size correlates with performance, but not perfectly. Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure˜4). However, this relationship is non-monotonic: moving from small to mid-sized models brings clear benefits, with notable exceptions such as Llama3-8B, which underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B. Among large models ($\approx$70B–120B+ parameters), performance gains diminish. Moreover, mid-sized models can match or outperform larger systems; for example, gpt-oss-20B clusters with Qwen-72B and Mistral-Large-2, while gpt-oss-120B approaches GPT-4/5 on the fine-grained benchmark. These results suggest that factors beyond parameter count influence model performance.
@@
    Figure 4: Model size vs. performance on high-level and fine-grained classification in Migrant (T1). Each model is plotted with its best-performing configuration (zero- or few-shot). The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
  • Examples help mainly mid-sized models. Most models benefit from few-shot prompting, with the most consistent improvements observed for mid-sized models such as Gemma models, Mistral-2-Large, Llama-3.3-70B and Llama-4-Scout. Smaller models like Llama-3-8B show only marginal gains, while GPT-4 demonstrates minimal improvement, likely due to its strong zero-shot performance. Overall, the effect of few-shot prompting is widespread but not uniform across model scales.
3.1.2 Task and label-related challenges
  • All models (and human annotators) perform worse on the fine-grained task, which indicates its complexity.
  • Solidarity is easier to detect than anti-solidarity or mixed stance (Figure˜5). Even smaller models perform well on solidarity, especially in the Woman category (e.g., BERT 0.77, Gemma-9B 0.70). In contrast, anti-solidarity and mixed stances remain challenging, with anti-solidarity being most difficult for Woman. This pattern is likely due in part to class imbalance (Table˜6 in the Appendix), especially because anti-solidarity is very rare for Woman and mixed is less common in both target groups. That said, GPT models (both closed-source models and gpt-oss-120B) achieve the strongest anti-solidarity performance (0.66–0.70 for Migrant; 0.53–0.65 for Woman), while mid-sized models reach 0.56–0.59. Only GPT-4/5, gpt-oss-120B and Llama-3.3-70B consistently capture mixed stances, with F1 scores between 0.30 and 0.50. Llama-4-Scout shows some asymmetry for Migrant: strong performance on solidarity (0.78) and anti-solidarity (0.63) but near failure on mixed stance (0.08).
  • On the fine-grained level, GPT-5, gpt-oss-120B, Llama-3.3-70B and Llama-4-Scout offer the most balanced performance on the fine-grained level (see Table˜7 in the Appendix). Best fine-grained macro F1 is achieved by GPT-5. Among non-GPT-4/5 models, Llama-3.3-70B and Llama-4-Scout follow. Nevertheless, all models perform poorly on specific subtypes, including empathic solidarity (e.g., 0.11 for Migrant with Llama-3.3-70B) and compassionate anti-solidarity (e.g., 0.06 for Migrant with gpt-oss-120B). Empathic anti-solidarity fails across all models. For Woman, all models consistently struggle with fine-grained anti-solidarity.